OpenStudy (anonymous):

Verify that the equation is an identity ; sin(45+ theta) - sin(45- theta) = square root 2 sin theta

OpenStudy (anonymous):

\[\sin(45+\theta)-\sin(45-\theta)=\sqrt{2}\sin\]

OpenStudy (midhun.madhu1987):